MULTIPLE CHOICE. Choose the n a tive that best completes the statement or answer the questi 1) Write the equilibri ne the equ Just give me the answers
Answer #1

1. The equilibrium constant expression for the reaction

2CH(g) + 302(g) = 2CO(g) + 4H2O(g) is

C) Ke= COH2014 (CH4)2 [0213

------------------------

2. Answer:- B) 0.324 mol

The reaction is H2(g) + Br2(g) = 2H Brg)

Initial mol of H2 = 0.682

mole of H2 at equilibrium = 0.566 mol

ie. mol of H2 reacted = (0.682 - 0.566) = 0.116 mol

Initial moles of Br2 = 0.440 mol

from reaction it is clear that H2 and Br2 reacts in 1:1 ratio.

therefore, moles of Br2 reacted = 0.116 mol

and

equilibrium moles of Br2 = (0.440 - 0.116) = 0.324 mol

-------------------

3) Answer :- C) at equilibrium

The given reaction 2.5O2(g) + O2(g) = 2.503(g)

Kc = 5

SO2) = 1.2M   02] = 0.45M  SO3 = 1.8M  

Now,

(S0312 Qc = [S02][02] (1.8) (1.2)(0.45)

i.e Qc = Kc ie. systemis at equilibrium.

--------------------

4) Answer :- B) Kp> Kc

The given reaction is

2CH(g) + 302(g) = 2CO(g) + 4H2O(g)

We know K, = K(RT)In

now, An = (molesof product – molesofreactant)

from reaction An = ((2+4) - (2+3) = 1

therfore, K, = K (RT) = K (RT)

ie. Kp > Kc

-------------------

5) Answer :- D) 0.033

----------

6) Answer:- C) KOK